We have been asked to do the following problem, for some reason it's just not sinking in. Any help that you could provide would be greatly appreciated.
One solution of 3x^2-7x+c=0 is 1/3. Find C and the other solution.
Thanks!
Lisa

Answer by venugopalramana(3286)

One solution of 3x^2-7x+c=0 is 1/3.
LET THE OTHER ROOT BE A...IN A QUADRATIC EQN.AX^2+BX+C=0..SUM OF ROOTS =-B/A AND PRODUCT OF ROOTS = C/A...HENCE.....
SUM OF ROOTS =A+1/3=-(-7)/3=7/3....SO A=7/3-1/3=6/3=2
PRODUCT OF ROOTS =A*1/3=2*1/3=2/3=C/3....SO C=2